You have just completed an analysis in which your degrees of freedom for all of your effects (main and interction) are the same. What type of analysis did you run?

Select one:

a. 2 x 2 x 2 between-subjects ANOVA

b. 1 x 4 between-subjects ANOVA

c. 2 x 1 between-subjects ANOVA

d. 3 x 3 between-subjects ANOVA
Answer #1

Solution:-

a. 2 x 2 x 2 between-subjects ANOVA

DFfactorA = a -1 = 2 -1 = 1

DFfactor B = b -1 = 2 -1 = 1

DFinteraction = (a - 1)(b - 1)= (2 -1)(2 - 1) = 1

Hence DFfactorA = DFfactor B = DFinteraction